Implicit Differentiation

Find the 2nd derivative of the following:

1) x=3+√x^2 + y^2

2) (x^2 + y^2)^3 = 8x^2y^2

3) x^2(y-x)^3 = 9


1
Expert's answer
2021-05-07T09:01:22-0400

1)


"x=3+\sqrt{x^2+y^2}"


"1=\frac{1}{2\sqrt{x^2+y^2}}(2x+2yy')"


"y'=\frac{\sqrt{x^2+y^2}-x}{y}"


"yy'=\sqrt{x^2+y^2}-x"


"(y')^2+yy''=\frac{x+yy'}{\sqrt{x^2+y^2}}-1"

from two previous equations


"(y')^2+yy''=0"


"y''=-\frac{(y')^2}{y}=-\frac{(\sqrt{x^2+y^2}-x)^2}{y^3}"


2)


"(x^2+y^2)^3=8x^2y^2"


"3(x^2+y^2)^2(2x+2yy')=16xy^2+16x^2yy'"


"6(x^2+y^2)(2x+2yy')^2+3(x^2+y^2)^2(2+2(y')^2+2yy'')="


"=16y^2+16x*2yy'+32xyy'+16x^2(y')^2+16x^2yy''"



"3x(x^2+y^2)^2-8xy^2=(8x^2y-3y(x^2+y^2)^2)y'"


"y'=\frac{3x(x^2+y^2)^2-8xy^2}{8x^2y-3y(x^2+y^2)^2}"


3)


"x^2(y-x)^3=9"


"2x(y-x)^3+x^2*3(y-x)^2(y'-1)=0"


"y'-1=\frac{-2(y-x)}{3x}=-\frac{2y}{3x}+\frac{2}{3}"


"y'=-\frac{2y}{3x}+\frac{5}{3}"


"y''=-\frac{2y'*3x-2y*3}{9x^2}=\frac{2y-2xy'}{3x^2}="


"=\frac{2y}{3x^2}-\frac{2}{3x}(-\frac{2y}{3x}+\frac{5}{3})="


"=\frac{10y}{9x^2}-\frac{10}{9x}=\frac{10(y-x)}{9x^2}"